PS-IS-121/RC

 

MINUTES OF THE PS-ISOLDE COORDINATION MEETING

held on 10 September 2002

 

 

Present:

B. Crepieux, R. Catherall, V. Fedosseev, G.J. Focker, T. Giles, A. Jokinen, G. Le Scornet, J. Lettry, M. Lindroos, S. Marzari, S. Meunier, A. Muller, J. Parra-Lopez, F. Wenander.

 

ISOLDE at URL: http://isolde.cern.ch

Consolidation projects at URL:

http://isolde-consolidation.web.cern.ch/isolde-consolidation/

PSMB meetings at URL:

http://cern.web.cern.ch/CERN/Divisions/PS/Meetings/PSMB.html

 

1. Separators:

      On Friday 6/9/02 at 05:15, the ISOLDE machine was the victim of a CERN-wide electrical breakdown. Most of the day was spent trying to re-start the machine. This included; starting vacuum sections, re-booting servers and magnet power supplies, and starting up the Nb target on the HRS. The experiment was able to continue in the evening.

 

2. Robots:

      The GPS robot is now working correctly. All the targets on the GPS side were removed with a minimum radioactive dose (479 uSv for 7 people). R. Catherall and Staubli were then able to identify the problem and correct it. One of the pick and place parameters had been altered causing the robot to drag the target rather than release it.

 

     

3. Vacuum

During the re-starting of the separators last friday, the empty emergency water reservoir prevented the powering of the target transformers. Normally, this reservoir should fill automatically after an emergency stop. While the vacuum group is looking in to this problem, a temporary method for filling the tank has been provided should this problem re-occur.

 

4. Beam lines

G. J. Focker reports that the FNBS is now working again.

 

5. Radiation Protection

The incorrect use of access cards both for the HRS secondary zone and the HT zone causes unnecessary delays. If one is not sure how to close a secondary door correctly please ask somebody who knows.

RP would like to remind all those handling radioactivity to wear suitable protective clothing.

 

6. Tunnel

A blocked dehumidifier caused a water leak in the tunnel. Although now repaired a full revision of this system will be planned for the shutdown.

  1. Muller reports that there is a radiation “hot spot” at the bending magnet 308 that deviates the beam to the HRS.

The quick coupling compressed air connectors of the shutter valve on the GPS front end had freed themselves. S. Marzari has re-connected them.

 

6. AOB

An inventory of all the keys to Building 179 and their owners will be made during the shut down period.

CERN management is preparing a position document that outlines the agreement between CERN and the ISOLDE collaboration for the operation of the REX-ISOLDE facility.

Rick Bartmann will visit ISOLDE from the 14th to 16th of October.

F. Wenander reports that he is currently testing a target with an ECR ion source and that a new ECR ion source is under development.

 

Absences

T. Giles from 13.09 to 07.10

 

 

Next meeting:

Tuesday 17 September 2002 at 11.00 in the EP Conference Room A, building 304-1-001.